Natural Gas Producers Inc. (NGPI) has an issue of 5-year, 9% annual coupon bonds outstanding. The bonds, which were originally issued 15 years ago, have a face value (FV) of $1,000, a yield-to-maturity (YTM) of 12%, and are noncallable. What is the current market price of NGPI’s bonds?
Answer #1

Coupon = 0.09 * 1000 = 90

Current price = coupon * [1 - 1 / (1 + r)nn] / r + FV / (1 + r)nn

Current price = 90 * [1 - / (1 + 0.12)5] / 0.12 + 1,000 / (1+0.12)5

Current price of bond = 90 * 3.604776 + 567.426856

Current price of bond = $891.86

Keys to use in a financial calculator

Fv 1000

PMT 90

I/Y 12

N 5

CPT PV